OpenStudy (anonymous):

Please help hard math will medal! Which pair of functions is not a pair of inverse functions? A. f(x) = x+1/6 and g(x) = 6x-1 B. f(x) = x-4/19 and g(x) = 19x+4 C. f(x) = x^5 and g(x) = 5(sqrt)x D. f(x) = x/x+20 and g(x) = 20/x-1 It's hard I'm thinking its D though?

OpenStudy (mathmale):

You could either find the inverse of each given function f(x), OR, alternatively, substitute function g(x) in each case for x within f(x) and determine whether the end result is x. If yes, the 2 functions are inverses of one another; otherwise, they are not.

OpenStudy (mathmale):

Example: "C. f(x) = x^5 and g(x) = 5(sqrt)x" In f(x)=x^5, replace x with 5sqrt(x): f(g(x)) = (5sqrt(x))^5 \[=5^5 * (\sqrt{x})^5\]

OpenStudy (mathmale):

Is this result = to x? If yes, f and g are inverse functions; if not, f and g are not inverses.

OpenStudy (mathmale):

Do the same thing with D. Replace "x" in f(x) with g(x)=20 / (x-1). Is the result x? or not x?

OpenStudy (jdoe0001):

so... do as mathmale indicated, and you'll see which one is it

OpenStudy (anonymous):

20/x ? So d?

OpenStudy (anonymous):

@mathmale

OpenStudy (jdoe0001):

what did you get for the inverse of D?

OpenStudy (mathmale):

I need to see how you've settled on option D. Which method did you use: finding the inverse of the first function, or substituting the 2nd function for "x" in the first function?

OpenStudy (anonymous):

I used math way lol this stuff is confusing

OpenStudy (anonymous):

So I'm guessing it's not d

OpenStudy (jdoe0001):

well... hmmm have you covered Inverse Functions yet? I mean, do you know how to get the inverse of a function? as mathmale stated above

OpenStudy (jdoe0001):

as far as mathway well, if your instructions for the exercise is use a calculator, and transcribe over whatever the calculator says then that's ok now, if you're supposed to do it, then I'd go with mathmale's way or even check your book, it should have how to get the inverse of a function

OpenStudy (anonymous):

It says NOT a pair, so would it be D?

OpenStudy (jdoe0001):

well... we dunno what the inverse of D is though

OpenStudy (jdoe0001):

so... get the inverse, see what you get maybe the one there IS the inverse, who knows

OpenStudy (anonymous):

D=-x

OpenStudy (jdoe0001):

well, the inverse of x/(x+20) is not -x so. recheck your inverse

OpenStudy (anonymous):

I made a typo on d It's 20x over x-1